D status 03 11 03 18
Download
1 / 10

DØ Status: 03/11 – 03/18 - PowerPoint PPT Presentation


  • 91 Views
  • Uploaded on

DØ Status: 03/11 – 03/18. Week integrated luminosity 1.1 pb -1 delivered (recorded by D0) 0.7 pb -1 lost due to Master Clock 1.0 pb -1 utilized (recorded runs) 65% Live Fraction… Excluding Master Clock issues Affected three stores last week (1068,1070,1072)

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about ' DØ Status: 03/11 – 03/18' - dorian-yang


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
D status 03 11 03 18
DØ Status: 03/11 – 03/18

  • Week integrated luminosity

    • 1.1 pb-1 delivered (recorded by D0)

      • 0.7 pb-1 lost due to Master Clock

    • 1.0 pb-1 utilized (recorded runs)

    • 65% Live Fraction…

    • Excluding Master Clock issues

      • Affected three stores

        last week (1068,1070,1072)

  • Data collection downtime

    • ~15 min lost Begin/ End Store

      • HV ramps, run config, L3/DAQ

    • Muon readout crates

      • 15% of total downtime

    • Remaining sources

      • L3/DAQ, CAL/CFT/SMT readout crates

L3 Input Rate

15 March 2002

~50 Hz

Muon

Store

1080

Store

1082

Events/15s

Hours

No store

VRC & SMT

Power Cycle

Alan L. Stone

Louisiana Tech University


Problems with master clock

Trigger framework problems began owl shift 3/12 (Store 1070)

Seq halt & clock parity errors

DØ Experts investigated

Potentially bad PCC module but no spare!

Called MCR for help

Steve Chappa – found spare PCC module

(but not certified)

Needed extension mods to provide sync for data taking

Successful TFW run

Using 2nd PCC module w/kluged phase lock loop circuit board from 1st PCC module

MC Timing output signals wrt beam x-ing has moved less than 2 ns compared to recent measurement (23-FEB-02)

We still don’t know why it failed…

Will have spare working PCC within the week

Problems with Master Clock

Phase Coherent Clock

3/12 13:00

3/13 10:00

3/13 13:30

Alan L. Stone

Louisiana Tech University


D detector status

We made no Controlled Access requests

Took advantage to make minor repairs & to do small detector maintenance jobs

Calorimeter: Full readout

T&C FPGA change moved timing by max 30 ns each readout crate

BLS power supplies stable

SMT: Full readout

No LV supply trips in past three weeks!

CFT: Steady improvement

All axials & 50% stereo boards

Total channel readout increases with each access & board swaps

Muon: Steady running

Readout crate errors are major issue  under study

29/500k events source of errors

DØ Detector Status

Sub-detector Report

Alan L. Stone

Louisiana Tech University

Alan L. Stone


Inter cryostat detector
Inter-Cryostat Detector

  • Joint effort: Univ. Texas-Arlington&Louisiana Tech Univ.

  • Enhances the hermeticity and uniformity of the Calorimeter System

    • Rapidly changing material profile & extra “dead” material between the Central & Endcap Calorimeters

    • ICD provides additional sampling in the ICR

  • Improves ET calculation & Jet Energy Resolution

    • Crucial role in coverage of 1.1 < || < 1.4

    • Reduce rate of fake ET

  • Commissioning of Calorimeter Readout

    • ICD channels “sprinkled” throughout most of the calorimeter readout system

Alan L. Stone

Louisiana Tech University


D detector
DØ Detector

ICD

Alan L. Stone

Louisiana Tech University


Inter cryostat region
Inter-Cryostat Region

 = 1.1

 = 1.4

Alan L. Stone

Louisiana Tech University


Stages of ICD installation from TDR to physics data taking

8

Alan L. Stone

Louisiana Tech University


Comparison of pre-Oct 2001 data and post-Nov 2001 data. The north part of the ICD (<0) was <5% instrumented prior to Oct 2001. The left plots (top & bottom) show the change in the average event energy in the north ICD channels pre/post shutdown. The right plots are of the south ICD, which has been fully instrumented since May 2001.

Pre-October

GeV

GeV

Post-November

GeV

GeV

North

South

Alan L. Stone

Louisiana Tech University


Icd status

50% in readout by end-May 2001 north part of the ICD (

Fully instrumented & part of global data taking since Nov 2001

Integrated into CAL Readout & Shift Guide

Stable HV system

Spare electronics & cables in stock

Working LED calibration system

Repair & maintenance during detector accesses

e.g. Replace Run 1 PMTs with new ones

Fine-tune energy scale & weights

Improve MC geometry & material representation

ICD Status

Done

To Do

Alan L. Stone

Louisiana Tech University


Plans for this week
Plans For This Week north part of the ICD (

  • Need 2 hr access (day & time TBD)

    • Muon VME crate (1 of 12) needs human intervention

  • Steady supply of small jobs … if offered additional access time

  • During studies Tuesday & Wednesday

    • Significant Event Server tests

    • Spare PCC tests

    • Muon readout

      • Test new code & new hardware

    • Test of multibuffering VRBCs

  • Global data taking during stores

    • Working on increasing data collection efficiency

Alan L. Stone

Louisiana Tech University


ad